OpenStudy (anonymous):

What is the remainder when 16,055 is divided by 16?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

1003.43

OpenStudy (anonymous):

16*1000 = 16000

OpenStudy (anonymous):

Medal pls

OpenStudy (anonymous):

so 55 = 16*3+r

OpenStudy (anonymous):

and 16055=16*1000+16*3+r= 16*(1000+3)+r

OpenStudy (anonymous):

so 55 is remainder

OpenStudy (anonymous):

no, the remainder must be less than 16 but greater than or equal to 0.

OpenStudy (anonymous):

r is the remainder

OpenStudy (anonymous):

1003

OpenStudy (anonymous):

16055 = 16*(1003) + r r is the remainder

OpenStudy (anonymous):

16*(1003) = 16048

OpenStudy (anonymous):

subtract and solve

OpenStudy (anonymous):

?

OpenStudy (anonymous):

16055 = 16048 + r solve for r, the remainder